Say there were 5 kids in class, the scores of the test were 90,85,95, 82 what is the other childs score when the average of the students is 75%.
Answer by checkley77(12844)

(90+85+95+82+x)/5=75
(352+x)/5=75 cross multiply.
352+x=5*75
352+x=375
x=375-352
x=23 ans. for the grade of the fifth student.
Proof:
(90+85+95+82+23)/5=75
375/5=75
375=5*75